OpenStudy (anonymous):

Solve the equation for x. Round your answer to two decimal places. 11-2 times log x=14 Can anyone help? Thanks!

OpenStudy (anonymous):

11-2log(x)=14 subtract 11 both sides -2log(x)=3 divide by -2 log(x)=-3/2 10^log(x)=10^(-3/2) x=10^(-3/2)